OpenStudy (anonymous):

z varies directly with x and inversely with y2. When x = 2 and y = 5, z = 8. What is the value of z when x = 4 and y = 9?

OpenStudy (mertsj):

\[z=\frac{kx}{y^2}\]

OpenStudy (mertsj):

Plug in the given values and find the constant, k.

OpenStudy (mertsj):

What did you get for k?

OpenStudy (anonymous):

Im looking at my notes but I cant figure out which number goes where? @Mertsj

OpenStudy (mertsj):

Why do you have to look at your notes? Why don't you read the problem? It clearly says x=2. That means replace x with 2

OpenStudy (mertsj):

Then it says y = 5. That means replace y with 5

OpenStudy (mertsj):

Then it says z = 8. That means replace z with 8

OpenStudy (anonymous):

ok, so I plugged in the numbers, 8=2k/5, and I made 2/5 a reciprocal and multiplied to 8 so it would even out right? and I got k= 1525.87, then I plugged that number in for the other side of the equation, z=1525.87(4)/9 and I got 678.16, which is not even close to the options I have.

OpenStudy (whpalmer4):

\[z = k\frac{x}{y^2}\]\[8= k\frac{2}{5^2}\]\[25*8 = 2k\]\[k=100\] You failed to square \(y\) in the denominator, plus some other mistake I'm not going to try to figure out.

OpenStudy (anonymous):

lol^,I got it already, thanks

OpenStudy (whpalmer4):

if you already have the answer, close the question so no one wastes time on it...